OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (anonymous):

@jim_thompson5910 @Mertsj @jdoe0001 @whpalmer4 @tkhunny @Preetha @Jamierox4ev3r @studyislandhelp @Awesome781

jimthompson5910 (jim_thompson5910):

f(x) = -4x + 7 f(g(x)) = -4(g(x)) + 7 ... replace EVERY x with g(x) f(g(x)) = -4(10x - 6) + 7 ... replace g(x) on the right side with 10x - 6 I'll let you finish up.

OpenStudy (jdoe0001):

\(\bf f(x) = -4x + 7 \qquad g(x) = {\color{red}{ 10x - 6}} \\ \quad \\ f(\quad g(x)\quad )=-4({\color{red}{ \square ? }}) + 7\)

OpenStudy (anonymous):

@jdoe0001 can you help me and elaborate on this?

OpenStudy (jdoe0001):

hmmm not much to elaborate really g(x) <-- the whole function, TAKES OVER any "x" in f(x) that's what f( g(x) ) means

OpenStudy (anonymous):

i dont understand how to work it out can you help me step by step

OpenStudy (jdoe0001):

\(\bf \bf f(x) = -4x + 7 \qquad g(x) = {\color{red}{ 10x - 6}} \\ \quad \\ f(\quad g(x)\quad )=-4({\color{red}{ 10x - 6}}) + 7\) <-- see it now?

OpenStudy (anonymous):

so it would be -40x+24

OpenStudy (jdoe0001):

yes, don't forget the +7 at the end

OpenStudy (jdoe0001):

\(\bf f(x) = -4x + 7 \qquad g(x) = {\color{red}{ 10x - 6}} \\ \quad \\ f(\quad g(x)\quad )=-4({\color{red}{ 10x - 6}}) + 7\implies -40x+24+7\implies -40x+31\)

OpenStudy (anonymous):

@jdoe0001 so the answer is b?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

can i ask you one more question @jdoe0001

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Let f(x) = x2 - 16. Find f-1(x). 4square root of x ±square root of x plus 16 x squared over 16 1 over quantity of x squared minus 16

OpenStudy (jdoe0001):

to find the "inverse relation" what we simply do is, swap about the variables and then solve for "y", so \(\bf f(x) = {\color{red}{ y}}={\color{blue}{ x}}^2 - 16\implies {\color{blue}{ x}}={\color{red}{ y}}^2 - 16\) and then you'd solve it for "y"

OpenStudy (anonymous):

would it be c? if not then its d

OpenStudy (jdoe0001):

\(\bf f(x) = {\color{red}{ y}}={\color{blue}{ x}}^2 - 16\implies {\color{blue}{ x}}={\color{red}{ y}}^2 - 16 \\ \quad \\ \textit{adding 16 to both sides }\implies x+16=y^2 \\ \quad \\ \textit{taking }\sqrt{\qquad }\textit{ to both sides}\implies \sqrt{x+16}=\sqrt{y^2} \\ \quad \\ \implies \sqrt{x+16}=y=f^{-1}(x)\)

OpenStudy (anonymous):

so its b?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks so much @jdoe0001 :)

OpenStudy (jdoe0001):

yw